Find the p-values when:

a. conducting a right-tail test, t=1.29 and n=45

b. conducting a two-tail test, t= -.91 and n=19

c. conducting a left-tail test, z= -2.02
Answer #1

a) Right-tail test, t=1.29 and n=45

df= n-1= 45-1 = 44

Distribution Plot T, df 44 0.3 0.2 0.1 0.10 19 0.0 0 1.29

P-value: 0.1019

b) t= -0.91 and n= 19

df= n-1= 18

Distribution Plot T, df 18 0.4- 0.3 0.2 0.1 0.1874 0.1874 0.0 0.91 00.91

P-value: 0.3748(Two-tailed)

c) Left tail: Z= -2.02

Distribution Plot Normal, Mean 0, StDev-1 0.4 0.3 E 0.2 0.1 0.02169 0.0 2.02

P-value: 0.02169